# Block encoding of matrix product operators

### v2

@@ Abstract
Quantum signal processing combined with quantum eigenvalue transformation has recently emerged as a unifying framework for several quantum algorithms. In its standard form, it consists of two separate routines: block encoding, which encodes a Hamiltonian in a larger unitary, and signal processing, which achieves an almost arbitrary polynomial transformation of such a Hamiltonian using rotation gates. The bottleneck of the entire operation is typically constituted by block encoding and, in recent years, several problem-specific techniques have been introduced to overcome this problem. Within this framework, we present a procedure to block-encode a Hamiltonian based on its matrix product operator (MPO) representation. More specifically, we encode every MPO tensor in a larger unitary of dimension $D+2$, where $D=\lceil\log(\chi)\rceil$ is the number of subsequently contracted qubits that scales logarithmically with the virtual bond dimension $\chi$. Given any system of size $L$, our method requires $L+D$ ancillary qubits in total, while the number of one- and two-qubit gates decomposing the block encoding circuit scales as $\mathcal{O}(L\cdot\chi^{2})$.
@@ I Introduction
In this process, the most significant bottleneck is represented by the block encoding of $\mathcal{H}$. It is well known that a unitary operator embedding $\mathcal{H}$ exists. However, decomposing such a unitary efficiently into one- and two-qubit gates requires some extra knowledge of the system and problem-specific techniques. In this context, we propose a versatile block encoding method based on a matrix product operator (MPO) representation of $\mathcal{H}$, where such an MPO form is typically known or can be constructed for systems of interest. Our method uses unitary dilations of the individual MPO tensors. Our approach achieves a block encoding with computational cost scaling only linearly with the system size $L$ and quadratically with the virtual bond dimension $\chi$ of the MPO, while the number of ancillary qubits scales like the sum of $L$ and $\log(\chi)$.
Our theoretical justification starts then in section II, where we provide a concise definition of MPOs within the context of tensor networks. This quick introduction is followed by the description of our algorithm in section III, which encodes such MPOs on a quantum computer. This task requires the use of ancillaries and measurements, as the MPO tensors are non-unitary in general.
We discuss the application of our method for QET in section IV and, specifically, we outline two methods to implement the signal processing operator: the original one described by Martyn et al. [3] and an alternative that requires no auxiliary qubits. The theoretical proof of the correctness of this method is provided in appendix A
We then perform a cost analysis in terms of two-qubit gates count, comparing our method to the linear combinations of unitaries (LCU) block encoding technique in section V. The results show that our technique is characterized by a scaling of $\mathcal{O}(L\cdot\chi^{2})$, while the cost of LCU varies depending on the number of unitaries composing the Hamiltonian.
Finally, we present some practical applications in section VI. We examine the MPO representation of the Ising and Heisenberg Hamiltonians and an exponentially decaying XY model, which exhibits a significant advantage compared to the use of LCU. We also consider a general tensor product of sums of Paulis as a straightforward example for realizing the entire QET circuit simulating an eigenstate filtering process.
@@ III Circuit representation
  • The encoded Hamiltonian $\mathcal{H}$ is normalized by a factor $\prod_{\ell=1}^{L}N_{\ell}$.
  • To retrieve $\mathcal{H}$, the dilation ancillaries must be all initialized and measured in the state |$\ket{0}$. The virtual bond ancillaries must also be prepared in state |$\ket{C}$ and measured in state |$\ket{R}$, representing the column and row vectors $C$ and $R$ previously defined. As a result, a post-selection operation is needed to fully implement $\mathcal{H}$ on a quantum circuit but, remarkably, the probability of measuring the right states is independent of the block encoding technique and number of ancillaries [10]:
@@
    $=\left\|\frac{\mathcal{H}}{\prod_{\ell=1}^{L}N_{\ell}}\ket{\psi_{\text{in}}}\right\|^{2}$
    $=\sum_{k}\left|\frac{\lambda_{k}}{\prod_{\ell=1}^{L}N_{\ell}}\right|^{2}\cdot\left|\braket{\lambda_{k}}{\psi_{\text{in}}}\right|^{2}$ (11)
    where |$\ket{\psi_{\text{in}}}$ is the input state on the physical wires and we have used the eigendecomposition of $H$: $H=\sum_{k}\lambda_{k}\ket{\lambda_{k}}\bra{\lambda_{k}}$ |. Interestingly enough, the probability is maximal when |$\ket{\psi_{\text{in}}}$ has a big overlap with the eigenvector corresponding to the maximal eigenvalue. Note that the normalization factors will, in general, affect the success probability. This aspect will be further discussed in sections  IV and  VII.
  • The cost for decomposing each $U_{A^{(\ell)}}$ into one- and two-qubit gates drastically varies for different applications and different native gate sets of the specific hardware. Nonetheless, in the most general case, it grows exponentially with the number of wires involved in each gate, equal to $D+2$, so that the number of two-qubit gates scales as $\mathcal{O}(L\cdot 4^{D})$ [11, 12]. Returning to the virtual bond dimension formalism and recalling Eq. (2), we can conclude that the computational cost for the MPO circuit scales linearly with the system size and quadratically with the virtual bond dimension as $\mathcal{O}(L\cdot\chi^{2})$.
@@ IV Quantum eigenvalue transformation with MPO block encoding
$\Pi_{\phi_{1}}U\prod_{k=2}^{(d-1)/2}\left(\Pi_{\phi_{2k}}U^{\dagger}\Pi_{\phi_{2k+1}}U\right)=\begin{pmatrix}\text{Poly}_{d}(\mathcal{H})&*\\ *&*\end{pmatrix}.$ (14)
@@
In section III, we have shown how the probability of success in the post-selection process doesn’t depend on the block encoding technique or the number of ancillaries. A similar analysis can be done for the full QET circuit:
@@
$=\left\|\text{Poly}_{d}\left(\frac{\mathcal{H}}{\prod_{\ell=1}^{L}N_{\ell}}\right)\ket{\psi_{\text{in}}}\right\|^{2}$
$=\sum_{k}\left|\text{Poly}_{d}\left(\frac{\lambda_{k}}{\prod_{\ell=1}^{L}N_{\ell}}\right)\right|^{2}\cdot\left|\braket{\lambda_{k}}{\psi_{\text{in}}}\right|^{2}$ (17)
Remarkably, this time, the probability of success is maximal when |$\ket{\psi_{\text{in}}}$ has a big overlap with the eigenvector |$\ket{\lambda_{k}}$ corresponding to the largest eigenvalue (in magnitude) after the transformation, and not of the original Hamiltonian. Note also that the normalization factors do not directly suppress the success probability anymore because they now appear in the argument of the polynomial transformation. Further discussion about this topic can be found in section VII.
@@ V Cost Analysis
Before illustrating some possible applications, we would like to discuss the cost of a general QET circuit built with our block encoding algorithm, in comparison to the linear combination of unitaries (LCU) technique [16, 2, 17] shown in Fig. 5. More specifically, we evaluate the computational cost based on the count of one- and two-qubit gates.
@@
For concreteness, we will consider the case where all the $U_{M}$ are Pauli strings. The total number of terms in the summation is denoted as $M$. In the worst case scenario, $M$ scales exponentially with the system size $L$ as $M=\mathcal{O}(4^{L})$. Still, there are important examples of polynomial or even linear scaling, like the Ising or Heisenberg Hamiltonian. We will first discuss the general case and then return to the specific examples later.
Block encoding ancillary qubits Block encoding computational cost (no extra auxiliaries) Martyn et al. [3] signal processing computational cost (1 extra auxiliary)
MPO Block Encoding $L+D$ $\mathcal{O}\left(L\cdot\chi^{2}\right)$ $\mathcal{O}(L+\log(\chi))$
LCU Block Encoding $\lceil\log(M)\rceil$ $\mathcal{O}\left(L\cdot M\cdot\log^{2}(M)\right)$ $\mathcal{O}(\log(M))$
LCU Block Encoding $M=\mathcal{O}(L)$ $\mathcal{O}\left(\log(L)\right)$ $\mathcal{O}\left(L^{2}\cdot\log^{2}(L)\right)$ $\mathcal{O}(\log(L))$
LCU Block Encoding $M\simeq 4^{L}$ $2L$ $\mathcal{O}\left(L^{3}\cdot 4^{L}\right)$ $\mathcal{O}(L)$
Table 1: Cost comparison between the LCU and MPO block encoding and signal processing in terms of number of single- and two-qubit gates and encoding ancillaries. The analysis assumes a cost for decomposing a general unitary on $n$ qubits of $\mathcal{O}(4^{n})$ [11, 12]. For the block encodings, we decompose multi-controlled Pauli gates with a quadratic cost without the use of extra auxiliary qubits [13, 18]. We consider Martyn et al.’s signal processing implementation [3] and a linear $n$-Toffoli decomposition which makes use of 1 extra auxiliary qubit [15]. The number of bond ancillary qubits $D$ scales logarithmically with respect to the MPO virtual bond dimension $\chi$ as $D=\lceil\log(\chi)\rceil$.
For LCU, one first adds $\lceil\log(M)\rceil$ ancillary qubits to the circuit that need to be prepared in the state:
|$\ket{\text{PREP}}=\sum_{m=1}^{M}\alpha_{m}\ket{m}$ (21)
with the same coefficients $\alpha_{m}$ as in Eq. (20). This preparation gate, as well as any other multi-qubit controlled gate that we will consider in this section, must be decomposed into one- and two-qubit gates. Generally speaking, this process has a cost that scales exponentially with the number of qubits [19], which means $\mathcal{O}(M)$.
The next step is known as the selection phase: here, each state |$\ket{m}$ controls one of the unitaries $U_{m}$. We have therefore $M$ Pauli strings acting on the $L$ system qubits and controlled by the $\lceil\log(M)\rceil$ ancillary qubits. Supposing that each Pauli gate can be implemented with a constant cost on a quantum computer, we get $\mathcal{O}(L)$ terms for each $U_{m}$. Then, for the decomposition of each multi-controlled Pauli, we have two possible choices:
@@
    One extra auxiliary qubit each, which allows for a linear scaling with the number of controls [14]. Note that the cited work refers to a $n$-controlled NOT, which can easily be converted into a multi-controlled $Z$ or $Y$ gate with constant cost. Assuming that the new extra qubit is entangled with the system and cannot be reused, in total this requires $L\cdot M$ extra auxiliaries for each $U$ and $U^{\dagger}$ in the QET circuit.
As a consequence, the cost of the decomposition for the full selection phase scales as $\mathcal{O}(L\cdot M\cdot\log^{\alpha}(M))$, where $\alpha$ is equal to $1$ or $2$ based on the presence of extra auxiliaries.
As a last step, we need to apply the inverse preparation gate on the ancillaries, with a final cost of $\mathcal{O}(2M+L\cdot M\cdot\log^{\alpha}(M))=\mathcal{O}(L\cdot M\cdot\log^{\alpha}(M))$.
@@
  • In the worst-case scenario, with $M$ scaling as $4^{L}$, we get $\mathcal{O}(L)$ encoding ancillary qubits and a total cost of $\mathcal{O}(L^{1+\alpha}\cdot 4^{L})$ gates. The case $\alpha=1$ would also require $\mathcal{O}(L\cdot 4^{L})$ extra auxiliaries.
  • For a general linear scaling, we get a total cost for the decomposition of $\mathcal{O}(L^{2}\cdot\log^{\alpha}(L))$ with only $\mathcal{O}(\log(L))$ encoding ancillaries and $\mathcal{O}(L^{2})$ optional extra auxiliaries for the decomposition for $\alpha=1$.
  • The aforementioned Ising and Heisenberg Hamiltonians also present a linear scaling, since they are defined as a linear sum of Pauli strings. However, because of the nearest-neighbor constraint, they can get even better complexity of $\mathcal{O}(L\cdot\log^{\alpha}(L))$, as the Pauli strings $U_{m}$ are only composed by one or two Pauli gates, and $\mathcal{O}(L)$ optional extra auxiliaries for the decomposition for $\alpha=1$.
We notice that adding extra auxiliaries in this framework brings a limited advantage compared to their scaling. For this reason, from now on we will only consider the case of no extra auxiliaries allowed for the decomposition of multi-controlled gates in the LCU block encoding.
@@
Table 1 summarizes the MPO and LCU block encoding costs, the latter with specific examples of $M$ scaling exponentially or linearly with $L$. We also added the signal processing decomposition cost, which scales only linearly with the number of ancillary qubits if we consider Martyn et al.’s implementation [3] and one extra auxiliary qubit [15].
Given the linear scaling with $L$ and the quadratic dependence on $\chi$, our block encoding turns out to be competitive compared to LCU. If we consider, for example, the aforementioned Ising and Heisenberg models, the corresponding MPO block encoding scales only linearly with $L$ since $\chi$ is constant, as it will be further discussed in section VI with other applications.
On the other hand, the relatively large number of ancillary qubits can make the signal processing phase more expensive. However, considering the entire QET circuit, we can conclude that this typically doesn’t constitute a limitation since the block encoding still prevails in the overall computational cost. The post-selection, moreover, is not affected by the number of ancillaries, as we have shown in Eq. (IV).
In conclusion, our method can constitute a valid alternative to LCU from a circuit size point of view. The two techniques are not directly comparable because $M$ and $\chi$ have different dependencies on the system size based on the application. We will show some example applications in section VI.

### v3

@@ Abstract
Quantum signal processing combined with quantum eigenvalue transformation has recently emerged as a unifying framework for several quantum algorithms. In its standard form, it consists of two separate routines: block encoding, which encodes a Hamiltonian in a larger unitary, and signal processing, which achieves an almost arbitrary polynomial transformation of such a Hamiltonian using rotation gates. The bottleneck of the entire operation is typically constituted by block encoding and, in recent years, several problem-specific techniques have been introduced to overcome this problem. Within this framework, we present a procedure to block-encode a Hamiltonian based on its matrix product operator (MPO) representation. More specifically, we encode every MPO tensor in a larger unitary of dimension $D+2$, where $D=\lceil\log(\chi)\rceil$ is the number of subsequently contracted qubits that scales logarithmically with the virtual bond dimension $\chi$. Given any system of size $L$, our method requires $L+D$ ancillary qubits in total, while the number of one- and two-qubit gates decomposing the block encoding circuit scales as $\mathcal{O}(L\cdot\chi^{2})$. Moreover, block encodings generally require post-selection. Our model exhibits a success probability decaying exponentially with $L$. Nevertheless, we present a way to mitigate this limitation for finite systems.
@@ I Introduction
In this process, the most significant bottleneck is represented by the block encoding of $\mathcal{H}$. It is well known that a unitary operator embedding $\mathcal{H}$ exists. However, decomposing such a unitary efficiently into one- and two-qubit gates requires some extra knowledge of the system and problem-specific techniques. In this context, we propose a versatile block encoding method based on a matrix product operator (MPO) representation of $\mathcal{H}$, where such an MPO form is typically known or can be constructed for systems of interest. Our method uses unitary dilations of the individual MPO tensors. Our approach achieves a block encoding with a number of one- and two-qubit gates scaling only linearly with the system size $L$ and quadratically with the virtual bond dimension $\chi$ of the MPO, while the number of ancillary qubits scales like the sum of $L$ and $\log(\chi)$.
The post-selection cost must also be accounted for in order to get a full picture of the computational cost. All block encoding methods share this issue, and it poses a challenge to the applicability of our algorithm, in particular, due to exponential decay in the system size. We will discuss this aspect in detail throughout the text and provide some ideas on how to overcome this limitation for finite systems.
Our theoretical justification starts then in section II, where we provide a concise definition of MPOs within the context of tensor networks. This quick introduction is followed by the description of our algorithm in section III, which encodes such MPOs on a quantum computer. This task requires the use of ancillaries, measurements and post-selection, as the MPO tensors are non-unitary in general.
We discuss the application of our method for QET in section IV and, specifically, we outline two methods to implement the signal processing operator: the original one described by Martyn et al. [3] and an alternative that requires no auxiliary qubits. The theoretical proof of the correctness of this method is provided in appendix A.
We then perform a cost analysis in terms of one- and two-qubit gates count, comparing our method to the linear combinations of unitaries (LCU) block encoding technique in section V. The results show that our technique is characterized by a scaling of $\mathcal{O}(L\cdot\chi^{2})$, while the cost of LCU varies depending on the number of unitaries composing the Hamiltonian. We also compare, on a theoretical level, the post-selection cost for both MPO and LCU block encodings, while a numerical analysis for systems of interest can be found in appendix B.
Finally, we present some practical applications in section VI. We examine the MPO representation of the Ising and Heisenberg Hamiltonians and an exponentially decaying XY model, which exhibits a significant advantage in gate decomposition cost compared to the use of LCU. We also consider a spinless one-dimensional Fermi-Hubbard model as an example of a fermionic system. At last, we construct a general tensor product of sums of Paulis as a straightforward use case for realizing the entire QET circuit simulating an eigenstate filtering process.
@@ III Circuit representation
  • The encoded Hamiltonian $\mathcal{H}$ is rescaled by a factor:
    $N_{\text{MPO}}=\prod_{\ell=1}^{L}N_{\ell},$ (11)
    Note that the normalization factors $N_{\ell}$ are typically bigger than $1$, which means that $N_{\text{MPO}}$ grows exponentially with the system size.
  • To retrieve $\mathcal{H}$, the dilation ancillaries must be all initialized and measured in the state |$\ket{0}$. The virtual bond ancillaries must also be prepared in state |$\ket{C}$ and measured in state |$\ket{R}$, representing the column and row vectors $C$ and $R$ previously defined. As a result, a post-selection operation is needed to fully implement $\mathcal{H}$ on a quantum circuit but, remarkably, the probability of measuring the right states is independent of the number of ancillaries [10]:
@@
    $=\left\|\frac{\mathcal{H}}{N_{\text{MPO}}}\ket{\psi_{\text{in}}}\right\|^{2}$
    $=\sum_{k}\left|\frac{\lambda_{k}}{N_{\text{MPO}}}\right|^{2}\cdot\left|\braket{\lambda_{k}}{\psi_{\text{in}}}\right|^{2}$ (12)
    where |$\ket{\psi_{\text{in}}}$ is the input state on the physical wires, we have used the eigendecomposition of $\mathcal{H}$: $\mathcal{H}=\sum_{k}\lambda_{k}\ket{\lambda_{k}}\bra{\lambda_{k}}$ | and the rescaling factor was defined in Eq. (11). Interestingly enough, the probability is maximal when |$\ket{\psi_{\text{in}}}$ has a big overlap with the eigenvector corresponding to the maximal eigenvalue. However, the rescaling factor $N_{\text{MPO}}$ affects the success probability with an exponential damping in $L$. This aspect will be further discussed in the following sections.
  • The cost for decomposing each $U_{A^{(\ell)}}$ into one- and two-qubit gates drastically varies for different applications and different native gate sets of the specific hardware. Nonetheless, in the most general case, it grows exponentially with the number of wires involved in each gate, equal to $D+2$, so that the number of two-qubit gates scales as $\mathcal{O}(L\cdot 4^{D})$ [11, 12]. Returning to the virtual bond dimension formalism and recalling Eq. (2), we can conclude that the gate decomposition cost for the MPO circuit scales linearly with the system size and quadratically with the virtual bond dimension as $\mathcal{O}(L\cdot\chi^{2})$.
@@ IV Quantum eigenvalue transformation with MPO block encoding
$\Pi_{\phi_{1}}U\prod_{k=1}^{(d-1)/2}\left(\Pi_{\phi_{2k}}U^{\dagger}\Pi_{\phi_{2k+1}}U\right)=\begin{pmatrix}\text{Poly}_{d}(\mathcal{H})&*\\ *&*\end{pmatrix}.$ (15)
@@
In section III, we have shown how the probability of success in the post-selection process does not depend on the number of ancillaries. A similar analysis can be done for the full QET circuit:
@@
$=\left\|\text{Poly}_{d}\left(\frac{\mathcal{H}}{N_{\text{MPO}}}\right)\ket{\psi_{\text{in}}}\right\|^{2}$
$=\sum_{k}\left|\text{Poly}_{d}\left(\frac{\lambda_{k}}{N_{\text{MPO}}}\right)\right|^{2}\cdot\left|\braket{\lambda_{k}}{\psi_{\text{in}}}\right|^{2}$ (18)
Remarkably, this time, the probability of success is maximal when |$\ket{\psi_{\text{in}}}$ has a big overlap with the eigenvector |$\ket{\lambda_{k}}$ corresponding to the largest eigenvalue (in magnitude) after the transformation, and not of the original Hamiltonian. Note also that the rescaling factor $N_{\text{MPO}}$ does not directly suppress the success probability anymore because it now appears in the argument of the polynomial transformation. Still, the exponential suppression term can squeeze the original eigenvalues towards 00, so that we would need an exponentially increasing $d$ to distinguish the different $\lambda_{k}$.
@@ V Cost Analysis
Before illustrating some possible applications, we would like to discuss the cost of a general QET circuit built with our block encoding algorithm in comparison to the linear combination of unitaries (LCU) technique [16, 2, 17] shown in Fig. 5. More specifically, we evaluate the computational cost based on the count of one- and two-qubit gates, and we compare the post-selection success probabilities for the two methods.
@@
For concreteness and without loss of generality, we will consider the case where all the $U_{M}$ are Pauli strings and the $\alpha_{m}$ are real and positive. The total number of terms in the summation is denoted as $M$. In the worst case scenario, $M$ scales exponentially with the system size $L$ as $M=\mathcal{O}(4^{L})$. Still, there are important examples of polynomial or even linear scaling, like the Ising or Heisenberg Hamiltonian. We will first discuss the general case and then return to the specific examples later.
For LCU, one first adds $\lceil\log(M)\rceil$ ancillary qubits to the circuit that need to be prepared in the state [2]:
|$\ket{\text{PREP}}=\frac{1}{\sqrt{N_{\text{LCU}}}}\sum_{m=1}^{M}\sqrt{\alpha_{m}}\ket{m}$ (22)
with the same coefficients $\alpha_{m}$ as in Eq. (21) and the factor $N_{\text{LCU}}$ defined as:
$N_{\text{LCU}}=\|\alpha\|_{1}=\sum_{m=1}^{M}|\alpha_{m}|.$ (23)
The latter is needed to make sure that |$\ket{\text{PREP}}$ is normalized, and this results in a rescaling of the Hamiltonian by a factor $N_{\text{LCU}}$.
This preparation gate, as well as any other multi-qubit gate that we will consider in this section, must be decomposed into one- and two-qubit gates. Generally speaking, this process has a cost that scales exponentially with the number of qubits [18], which means $\mathcal{O}(M)$.
The next step is known as the selection phase: here, each state |$\ket{m}$ controls one of the unitaries $U_{m}$. We have then $M$ Pauli strings acting on the $L$ system qubits and controlled by the $\lceil\log(M)\rceil$ ancillary qubits. Supposing that each Pauli gate can be implemented with a constant cost on a quantum computer, we get $\mathcal{O}(L)$ terms for each $U_{m}$. Then, for the decomposition of each multi-controlled Pauli, we have two possible choices:
@@
    One dirty auxiliary qubit, which allows for a linear scaling with the number of controls [14] and can be reused by all the multi-controlled Paulis. Note that the cited work refers to a $n$-controlled NOT, which can easily be converted into a multi-controlled $Z$ or $Y$ gate with constant cost.
As a consequence, the cost of the decomposition for the full selection phase scales as $\mathcal{O}(L\cdot M\cdot\log^{\beta}(M))$, where $\beta$ is equal to $1$ or $2$ based on the presence of the extra dirty ancilla.
As a last step, we need to apply the inverse preparation gate on the ancillaries, with a final cost of $\mathcal{O}(2M+L\cdot M\cdot\log^{\beta}(M))=\mathcal{O}(L\cdot M\cdot\log^{\beta}(M))$.
@@
  • In the worst-case scenario, with $M$ scaling as $4^{L}$, we get $\mathcal{O}(L)$ ancillary qubits and a total cost of $\mathcal{O}(L^{1+\beta}\cdot 4^{L})$ gates.
  • For a general linear scaling, we get a total cost for the decomposition of $\mathcal{O}(L^{2}\cdot\log^{\beta}(L))$ with only $\mathcal{O}(\log(L))$ ancillaries.
  • The aforementioned Ising and Heisenberg Hamiltonians also present a linear scaling since they are defined as a linear sum of Pauli strings. However, because of the nearest-neighbor constraint, they can get even better complexity of $\mathcal{O}(L\cdot\log^{\beta}(L))$, as the Pauli strings $U_{m}$ are only composed by one or two Pauli gates.
@@
Finally, we consider the post-selection cost for both methods. In section III, and more specifically in Eq. (III), we have derived the post-processing success probability for the MPO block encoding. The same calculations bring us to a similar result for LCU:
$p_{\text{success}}=\sum_{k}\left|\frac{\lambda_{k}}{N_{\text{LCU}}}\right|^{2}\cdot|\braket{\lambda_{k}}{\psi_{\text{in}}}|^{2}$ (24)
Here, the damping factor $N_{\text{LCU}}$ depends on the sum of the terms $\alpha_{m}$, as shown in Eq. (23). We can then expect a linear scaling in $M$, compared to the exponential dependence on $L$ of the MPO case. Still, comparing these methods is not trivial, as the scaling of $M$ with respect to $L$ depends on the system considered. In appendix B, we provide numerical comparisons for the two methods. It’s also worth remembering that the success probability formulas change if we consider the entire QET circuit, as shown in Eq. (IV). This, of course, stays true also in the LCU case.
Table 1 summarizes the MPO and LCU block encoding decomposition costs, the latter with specific examples of $M$ scaling exponentially or linearly with $L$. We also added references to the post-selection of the Hamiltonians and the signal processing decomposition cost. The latter scales only linearly with the number of ancillary qubits if we consider Martyn et al.’s implementation [3] and one extra auxiliary qubit [14].
Block encoding ancillary qubits Block encoding computational cost $p_{\text{success}}$ of retrieving $\mathcal{H}$ Martyn et al. [3] signal processing computational cost (1 extra auxiliary)
MPO Block Encoding $L+D$ $\mathcal{O}\left(L\cdot\chi^{2}\right)$ Eq. (III) $\mathcal{O}(L+\log(\chi))$
LCU Block Encoding $\lceil\log(M)\rceil$ $\mathcal{O}\left(L\cdot M\cdot\log^{\beta}(M)\right)$ Eq. (24) $\mathcal{O}(\log(M))$
LCU Block Encoding $M=\mathcal{O}(L)$ $\mathcal{O}\left(\log(L)\right)$ $\mathcal{O}\left(L^{2}\cdot\log^{\beta}(L)\right)$ Eq. (24) $\mathcal{O}(\log(L))$
LCU Block Encoding $M\simeq 4^{L}$ $2L$ $\mathcal{O}\left(L^{1+\beta}\cdot 4^{L}\right)$ Eq. (24) $\mathcal{O}(L)$
Table 1: Comparison of the decomposition cost between the LCU and MPO block encoding and signal processing in terms of number of single- and two-qubit gates, of the encoding ancillaries’ number and of the post-selection’s success probability (without QET). The analysis assumes a cost for decomposing a general unitary on $n$ qubits of $\mathcal{O}(4^{n})$ [11, 12]. For the LCU’s select phase, we can decompose multi-controlled Pauli gates with a quadratic cost without the use of extra auxiliary qubits [13, 19] and $\beta=2$, or with one dirty ancilla [14] and $\beta=1$. We consider Martyn et al.’s signal processing implementation [3] and the linear $n$-Toffoli decomposition making use of 1 extra auxiliary qubit [14]. The number of bond ancillary qubits $D$ scales logarithmically with respect to the MPO virtual bond dimension $\chi$ as $D=\lceil\log(\chi)\rceil$.
In summary, given the linear scaling with $L$ and the quadratic dependence on $\chi$, our block encoding turns out to be competitive compared to LCU when considering the circuit size. If we consider, for example, the aforementioned Ising and Heisenberg models, the corresponding MPO block encoding scales only linearly with $L$ since $\chi$ is constant, as it will be further discussed in section VI with other applications.
On the other hand, post-processing can present a lower success probability in the asymptotic limit. The relatively large number of ancillary qubits can also make the signal processing phase more expensive, but considering the entire QET circuit, we can conclude that this doesn’t constitute a limitation since the block encoding still prevails in the overall computational cost.
Finally, we would like to remark that the two techniques are not directly comparable because $M$ and $\chi$ have different dependencies on the system size based on the application. We will show some example applications in section VI.